OpenStudy (anonymous):

(topic: sol. of triangle) In a triangle ABC if a^2+b^2+c^2=ac+ab sqrt(3) then prove it is right angled. (a,b,c are sides opp. A,B,C)

hartnn (hartnn):

a^2+b^2+c^2=ac+2ab sqrt(3) /2 a^2+b^2+c^2=ac+2ab cos 30 so we can assume angle C = 30 , then by cos law c^2=a^2+b^2-2ab cos 30 , so , 2ab cos 30 = a^2+b^2 -c^2 substituting this in a^2+b^2+c^2=ac+2ab cos 30 we get, 2c^2 = ac ----> c/a = 1/2 Then from sine Law, sin A/ sin C = c/a gives us A =90.

hartnn (hartnn):

neither do i know whether its correct :P just gave it a try..... if i had known for sure, i would make him reach the answer...

OpenStudy (phi):

It's not hard to show that if you have a 30-60-90 triangle, then a^2+b^2+c^2 = ac+bc sqrt(3) where c is the hypotenuse. But going the other way is more difficult. I can massage things to \[ a \cos B + b \cos A -\frac{ab}{c} \cos(A+B)= a\cdot \frac{\sqrt{3}}{2} + b \cdot \frac{1}{2} \] then argue that the coefficients of a and b must match up. Is that a proof?